# Symmetry-Constrained Generation of Diverse Low-Bandgap Molecules with Monte Carlo Tree Search

### v1

@@ 1 Introduction
Over the past half-decade, generative models have received tremendous attention in molecular design (Schwalbe-Koda & Gómez-Bombarelli, 2020; Elton et al., 2019), using a variety of algorithms and molecular representations (Segler et al., 2018; Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Olivecrona et al., 2017; Mercado et al., 2021; Jin et al., 2018; Flam-Shepherd et al., 2021). However, without special attention to synthetic accessibility, off-the-shelf models often generate non-synthesizable designs (Gao & Coley, 2020). Several existing approaches incorporate knowledge about synthesizability into generative models, either implicitly or explicitly.
@@ 2.1 Fragment Identity, Connectivity and Symmetry
Our previous work developed an automated pipeline to mine USPTO patents and create domain-focused unsupervised datasets of molecular structures (Subramanian et al., 2023). These patent-derived molecules likely possess two important characteristics: high performance (justifying the investment in a patent application) and synthetic accessibility. Consequently, the fragments present in these structures, along with their reactive positions (points of connectivity), potentially correlate with both synthetic accessibility and domain-specific properties. Moreover, the symmetry of reactive positions relative to each other encodes information about the synthetic pathways used to create these molecules. To effectively leverage this structural and synthetic information, we have developed a two-stage method: Fragment Decomposition: We designed an algorithm that extracts fragments and their reactive positions from patent-mined structures, while marking each reactive position in a way that faithfully represents the identity of the substructure attached to it. This preserves important contextual information about molecular connectivity. Constrained Generation: We implemented a generation algorithm based on MCTS that constrains the output to molecules composed solely of patent-extracted fragments. These fragments are attached at their corresponding reactive positions while adhering to the symmetry requirements observed in the original patent structures. This approach aims to maintain synthetic feasibility while exploring novel molecular configurations. By combining these two components, we aim to create a framework for generating molecules that are not only likely to be synthetically accessible but also tailored to the specific property requirements of organic electronic materials. This method leverages the information embedded in patent-derived structures to guide the exploration of chemical space towards promising and realizable candidates.
@@ 2.1.1 Fragment Decomposition
In step 1, the reaction SMARTS is applied to the molecule to get all possible fragments that are obtainable from deletion of one bond in the molecule. The vocabulary of unique fragments is created, and each is assigned a unique ID. For the IDs, we use numbers $[0,N)$ where $N$ is the number of fragments in the vocabulary. Details on how this is actually represented in our code implementation are given in SI Section E.
@@ 2.1.2 Generation with MCTS
We frame the problem of designing a molecule as a series of building block/molecular fragment choices. MCTS is an efficient method of navigating large exploration spaces, that learns to optimize a chosen reward function by iteratively updating value estimates for action choices in the tree (Kocsis & Szepesvári, 2006; Coulom, 2006; Świechowski et al., 2023; Silver et al., 2016). Since our objective is to design low-bandgap molecules, our reward function of choice is the negative bandgap as predicted by Chemprop (Heid et al., 2023), a message-passing neural network (MPNN) architecture that we trained on labels obtained from TD-DFT calculations. Each iteration of the MCTS algorithm consists of four sequential steps: selection, expansion, rollout, and backpropagation. The selection step starts with the root node and traverses the already explored portions of the tree to reach a leaf node. In our work, the path from the root node to the leaf node represents the construction of a molecular structure, where each node corresponds to a partial or complete structure. We follow the standard Upper Confidence Bound applied to Trees (UCT) policy (Kocsis & Szepesvári, 2006) to choose an action at each timestep of traversal.
@@ 2.4.2 Active Learning for Data Collection
Figure 4: Final generated candidates and their TD-DFT bandgaps K-means clustering was performed on final 100 molecules from each category into 5 clusters (based on Morgan fingerprints). The molecules shown are the lowest-bandgap molecules chosen from each cluster. (a) Y6-derivatives MDP, (b) Patent-extracted fragments MDP. Bandgap of the last molecule is denoted with asterisk (*) because the geometry had to be fixed before the TD-DFT calculation was performed. More details are provided in Section D of SI.
@@ 3 Results and Discussion
Design strategies such as spatially separating donating and accepting fragments for a “push-pull” effect are also known to have a lowering effect on bandgaps (Pirotte et al., 2018). Among the generated patent-derivatives, this design pattern can be seen in the first and fourth molecules of subfigure (b) which contain electron-rich cores and electron-withdrawing end-groups. We additionally observe the presence of highly reactive sites in some of these structures (such as hydrogen atoms on aromatic amines of second, fourth, and fifth molecules, and chlorine atoms on the fourth molecule). We replace the reactive positions with methyl groups and perform TD-DFT calculations on the “corrected” structures to verify that optical properties do not drastically change even after the substitution. These results can be found in Section D of SI.
Finally, the TD-DFT bandgap calculations that we report work under the assumption of 0 K temperature. Furthermore, a more realistic device-like environment would contain thin-films containing several molecules packed together. We perform coupled MD/TD-DFT calculations on the final 10 molecules to better represent finite temperature and molecular aggregation effects. We report absorption spectra calculated with these methods, and more details about this methodology in section C of the SI.
@@ 5 Acknowledgements
Appendix A Preprocessing of Patent-Mined Fragments
@@
Appendix B Correlation Between DFT and Experimental Bandgaps
We performed TD-DFT calculations on a subset of experimental molecules from the list shown in Figure 7 for which we had access to experimentally measured bandgaps. We see $R^{2}$ value of 0.57, and a Spearman rank correlation coefficient (SRCC) of 0.74. While the magnitudes are not well calibrated, the trends of experiments are well-captured by TD-DFT, making minimization a suitable objective for MCTS as opposed to optimization for targeted bandgap values. As a future extension of this work, it could be possible to fit a calibration function between TD-DFT and experimental bandgaps and use the calibrated rewards for training. While it could be somewhat reasonable to extrapolate such calibrations for the Y6 MDP, it might be more challenging to obtain representative experimental examples for the patents MDP on which a calibration fit can be performed.
@@ Appendix B Correlation Between DFT and Experimental Bandgaps
Figure 5: Correlation between TD-DFT and experimental bandgaps $R^{2}$ score and Spearman rank correlation coefficient are shown on top left.
Appendix C Finite Temperature and Aggregation Effects with Coupled MD/TD-DFT Calculations
C.1 Computational Methods
Two of the patent-derived molecules (the first and fifth molecules in Figure 4) contain an isocyanide moiety, which could not be parametrized correctly using the LigParGen server. Therefore, for these two molecules, we substituted the isocyanide (–NC) functional group with cyanide (–CN) for this analysis. This substitution does not impact the band gap calculations significantly, with differences in bandgaps for the optimized geometries of only 2 and 6 meV, respectively, indicating negligible effects on the result analysis.
C.2 Absorption Spectra of Final Candidates
Appendix D Structural Modifications
For the fifth molecule shown in Figure 4, we identified that errors in atomic connectivity were introduced during the xTB optimization step. We fixed the geometry by bypassing the xTB optimization step altogether. We instead directly performed BP86 DFT optimization on the five lowest energy conformers obtained from RDKit ETKDG algorithm, and performed the TD-DFT calculation on the lowest energy BP86 conformer.
As mentioned in Section 3, we performed TD-DFT calculations on methyl-substituted versions of reactive molecules to confirm that the bandgap does not change significantly after substitution. Bandgaps of second, fourth, and fifth molecules changed very minimally from 0.998, 1.09, and 1.470, to 0.978, 1.049, and 1.470 eV respectively.
Appendix E Reactive Positions in Practice
While we explain the algorithm in Section 2.1.1 using IDs belonging to $[0,N)$, in the implementation, we use Helium isotopes with IDs ranging from $[100,100+N)$ where $N$ is the number of fragments in the vocabulary, to preserve reactive position identities in the RDKit Mol object. The range was chosen arbitrarily but to be large enough so that there are no clashes in atomic mass with other elements in the dataset.
Appendix F Fragments in Y6 MDP and Experimental Comparisons

### v2

@@ 1 Introduction
Over the past half-decade, generative models have received tremendous attention in molecular design (Subramanian et al., 2024; Schwalbe-Koda & Gómez-Bombarelli, 2020; Elton et al., 2019), using a variety of algorithms and molecular representations (Segler et al., 2018; Bjerrum & Threlfall, 2017; Gómez-Bombarelli et al., 2018; Kusner et al., 2017; Olivecrona et al., 2017; Mercado et al., 2021; Jin et al., 2018; Flam-Shepherd et al., 2021). However, without special attention to synthetic accessibility, off-the-shelf models often generate non-synthesizable designs (Gao & Coley, 2020). Several existing approaches incorporate knowledge about synthesizability into generative models, either implicitly or explicitly.
@@ 2.1 Fragment Identity, Connectivity and Symmetry
Our previous work developed an automated pipeline to mine USPTO patents and create domain-focused unsupervised datasets of molecular structures (Subramanian et al., 2023). These patent-derived molecules likely possess two important characteristics: high performance (justifying the investment in a patent application) and synthetic accessibility. Consequently, the fragments present in these structures, along with their reactive positions (points of connectivity), potentially correlate with both synthetic accessibility and domain-specific properties. Moreover, the symmetry of reactive positions relative to each other encodes information about the synthetic pathways used to create these molecules. To effectively leverage this structural and synthetic information, we have developed a two-stage method: Fragment Decomposition: We designed an algorithm that extracts fragments and their reactive positions from patent-mined structures, while marking each reactive position in a way that faithfully represents the identity of the substructure attached to it. This preserves important contextual information about molecular connectivity. It is this connectivity and identity-preserving property that we refer to as ”symmetry” throughout this paper. Constrained Generation: We implemented a generation algorithm based on MCTS that constrains the output to molecules composed solely of patent-extracted fragments. These fragments are attached at their corresponding reactive positions while adhering to the symmetry requirements observed in the original patent structures. This approach aims to maintain synthetic feasibility while exploring novel molecular configurations. By combining these two components, we aim to create a framework for generating molecules that are not only likely to be synthetically accessible but also tailored to the specific property requirements of organic electronic materials. This method leverages the information embedded in patent-derived structures to guide the exploration of chemical space towards promising and realizable candidates.
@@ 2.1.1 Fragment Decomposition
In step 1, the reaction SMARTS is applied to the molecule to get all possible fragments that are obtainable from deletion of one bond in the molecule. The vocabulary of unique fragments is created, and each is assigned a unique ID. For the IDs, we use numbers $[0,N)$ where $N$ is the number of fragments in the vocabulary. Details on how this is actually represented in our code implementation are given in SI Section F.
@@ 2.1.2 Generation with MCTS
We frame the problem of designing a molecule as a series of building block/molecular fragment choices. MCTS is an efficient method of navigating large exploration spaces, that learns to optimize a chosen reward function by iteratively updating value estimates for action choices in the tree (Kocsis & Szepesvári, 2006; Coulom, 2006; Świechowski et al., 2023; Silver et al., 2016). Since our objective is to design low-bandgap molecules, our reward function of choice is the negative bandgap as predicted by Chemprop (Heid et al., 2023), a message-passing neural network (MPNN) architecture that we trained on labels obtained from TD-DFT calculations. While more details about TD-DFT calculations are provided in Section 2.4.1, we note here that our use of the term ”bandgap” throughout this paper refers to the lowest-energy singlet vertical excitation energies. Each iteration of the MCTS algorithm consists of four sequential steps: selection, expansion, rollout, and backpropagation. The selection step starts with the root node and traverses the already explored portions of the tree to reach a leaf node. In our work, the path from the root node to the leaf node represents the construction of a molecular structure, where each node corresponds to a partial or complete structure. We follow the standard Upper Confidence Bound applied to Trees (UCT) policy (Kocsis & Szepesvári, 2006) to choose an action at each timestep of traversal.
@@ 2.4.2 Active Learning for Data Collection
Figure 4: Final generated candidates and their TD-DFT bandgaps K-means clustering was performed on final 100 molecules from each category into 5 clusters (based on Morgan fingerprints). The molecules shown are the lowest-bandgap molecules chosen from each cluster. (a) Y6-derivatives MDP, (b) Patent-extracted fragments MDP. Bandgap of the last molecule is denoted with asterisk (*) because the geometry had to be fixed before the TD-DFT calculation was performed. More details are provided in Section E of SI.
@@ 3 Results and Discussion
Design strategies such as spatially separating donating and accepting fragments for a “push-pull” effect are also known to have a lowering effect on bandgaps (Pirotte et al., 2018). Among the generated patent-derivatives, this design pattern can be seen in the first and fourth molecules of subfigure (b) which contain electron-rich cores and electron-withdrawing end-groups. We additionally observe the presence of highly reactive sites in some of these structures (such as hydrogen atoms on aromatic amines of second, fourth, and fifth molecules, and chlorine atoms on the fourth molecule). We replace the reactive positions with methyl groups and perform TD-DFT calculations on the “corrected” structures to verify that optical properties do not drastically change even after the substitution. These results can be found in Section E of SI.
Finally, the TD-DFT bandgap calculations that we report work under the assumption of 0 K temperature. Furthermore, a more realistic device-like environment would contain thin-films containing several molecules packed together. We perform coupled MD/TD-DFT calculations on the final 10 molecules to better represent finite temperature and molecular aggregation effects. We report absorption spectra calculated with these methods, and more details about this methodology in section B of the SI.
@@ 4 Conclusions
5 Data and Code Availability
All data and code accompanying this article, are available at https://github.com/learningmatter-mit/symmetry-mcts.
@@ 6 Acknowledgements
Appendix A Acquisition Function for AL
Expected improvement is an bayesian acquisition function that builds on top of probability of improvement. While a more detailed discussion can be seen in Kamperis (2021), we reproduce a brief excerpt of the derivation below.
A.1 Probability of Improvement (PI)
Given a function $f$, our task is to estimate the probability that $f(x)>f(x^{\star})$ where $x^{\star}$ is a previous optimum. If improvement is defined as,
$I(x)=max(0,f(x)-f(x^{\star})),$
If $f(x)$ is treated as a random variable following the gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$, then using the reparameterization trick, we can rewrite $I(x)$ as
$I(x)=max(0,\mu(x)+\sigma(x)z-f(x^{\star})),z\sim\mathcal{N}(0,1)$
Then the probability of improvement
$PI(x)=Pr(I(x)>0)=\Phi(\frac{\mu(x)-f(x^{\star})}{\sigma(x)}),$
where $\Phi(z)=CDF(z)$.
A.2 Expected Improvement (EI)
EI, unlike PI, is the expected value of improvement $I(x)$, which gives us an estimate of the magnitude of improvement rather than just the probability of improvement.
It is defined as
[…]
where $\varphi(z)=\frac{1}{\sqrt{2\pi}}exp(-z^{2}/2)$. The integral can be split as
[…]
The first term is zero since $I(x)=0$.
[…]
where $\Phi$ and $\varphi$ are the cumulative distribution function (CDF) and probability density function (PDF) of a normal distribution, respectively.
Appendix B Finite Temperature and Aggregation Effects with Coupled MD/TD-DFT Calculations
B.1 Computational Methods
Two of the patent-derived molecules (the first and fifth molecules in Figure 4 of the main text) contain an isocyanide moiety, which could not be parametrized correctly using the LigParGen server. Therefore, for these two molecules, we substituted the isocyanide (–NC) functional group with cyanide (–CN) for this analysis. This substitution does not impact the band gap calculations significantly, with differences in bandgaps for the optimized geometries of only 2 and 6 meV, respectively, indicating negligible effects on the result analysis.
B.2 Absorption Spectra of Final Candidates
Appendix C Preprocessing of Patent-Mined Fragments
@@
Appendix D Correlation Between DFT and Experimental Bandgaps
We performed TD-DFT calculations on a subset of experimental molecules from the list shown in Figure 6 for which we had access to experimentally measured bandgaps. The correlation plot between experimental and DFT values are shown in Figure 7. We see $R^{2}$ value of 0.57, and a Spearman rank correlation coefficient (SRCC) of 0.74. While the magnitudes are not well calibrated, the trends of experiments are well-captured by TD-DFT, making minimization a suitable objective for MCTS as opposed to optimization for targeted bandgap values. As a future extension of this work, it could be possible to fit a calibration function between TD-DFT and experimental bandgaps and use the calibrated rewards for training. While it could be somewhat reasonable to extrapolate such calibrations for the Y6 MDP, it might be more challenging to obtain representative experimental examples for the patents MDP on which a calibration fit can be performed.
@@ Appendix D Correlation Between DFT and Experimental Bandgaps
Figure 7: Correlation between TD-DFT and experimental bandgaps. $R^{2}$ score and Spearman rank correlation coefficient are shown on the top left.
Appendix E Structural Modifications
For the fifth molecule shown in Figure 4 of the main text, we identified that errors in atomic connectivity were introduced during the xTB optimization step. We fixed the geometry by bypassing the xTB optimization step altogether. We instead directly performed BP86 DFT optimization on the five lowest energy conformers obtained from RDKit ETKDG algorithm, and performed the TD-DFT calculation on the lowest energy BP86 conformer.
As mentioned in Section 3 of the main text, we performed TD-DFT calculations on methyl-substituted versions of reactive molecules to confirm that the bandgap does not change significantly after substitution. Bandgaps of second, fourth, and fifth molecules changed very minimally from 0.998, 1.09, and 1.470, to 0.978, 1.049, and 1.470 eV respectively.
Appendix F Reactive Positions in Practice
While we explain the algorithm in Section 2.1.1 of the main text using IDs belonging to $[0,N)$, in the implementation, we use Helium isotopes with IDs ranging from $[100,100+N)$ where $N$ is the number of fragments in the vocabulary, to preserve reactive position identities in the RDKit Mol object. The range was chosen arbitrarily but to be large enough so that there are no clashes in atomic mass with other elements in the dataset.
Appendix G Fragments in Y6 MDP
We illustrate the fragments along with their reactive positions in Figure 8.